OpenStudy (anonymous):

Find the length of the curve x=3y^(4/3)-(3)/(32)y^(2/3) , 8<=y<=21

OpenStudy (irishboy123):

just crank it out \[\int\limits ds = \int\limits_{y = 8}^{21} \sqrt{(\frac{dx}{dy})^2 + 1} \space{ }dy\]